You have $100,000 that you want to invest in a risky asset portfolio; M. Portfolio M has an expected return of 20% and a standard deviation of 30%. You also want to borrow an extra $50,000 at a risk-free rate of 8% and invest in the risky portfolio. Calculate the expected return and standard deviation of this portfolio. What type of portfolio is this and what type of investors will invest in this portfolio?
Answer #1

Expected return=1.5*20%-0.5*8%=26%

Standard Deviation=1.5*30%=45%

Leveraged Portfolio

Risk loving or risk neutral investors